OpenStudy (anonymous):

what is the average of 29 consecutive integers?

OpenStudy (anonymous):

the fifteenth

OpenStudy (anonymous):

how? and why?

OpenStudy (anonymous):

(29+1)/2 = 15

OpenStudy (anonymous):

let the first one =x, so they are x, x+1,x+2...x+28 ave=(x+x+1+x+2+..+x+28)/29 =(29x+(1+28)*29/2)/29 =x+14 it is the fifteenth one

OpenStudy (anonymous):

Sum of n consecutive natural numbers is n(n+1)/2